A pilot flew an average of 300 mph on a flight out. On the return flight to the same airport, he flew at an average speed of 500 mph. The total flight time was 8 hours. How far did he fly each way?
Answer by mananth(16946)

A pilot flew an average of 300 mph on a flight out. On the return flight to the same airport, he flew at an average speed of 500 mph. The total flight time was 8 hours. How far did he fly each way?
...
let distance be x miles
speed out = 300 mph
time = distance/speed
time out = x/300
...
similarly time return = x/500
..
time out + time return = 8 hours
x/300 + x/500 =8
(500x+300x)/500*300=8
800x= 8*500*300
/800
x= 8*500*300/800
x=1500 miles
...
m.ananth@hotmail.ca